# Deeper quantum circuits via pseudo-twirling coherent errors mitigation in non-Clifford gates

### v1

@@ I.2 Introducing the Pseudo Twirling Protocol
To introduce the concept of pseudo twirling we start with the following observation. Consider a simple non-Pauli gate, such as the two-qubit gate $U_{zz}(\theta)=e^{-i\theta\mathcal{H}_{zz}}$, where for simplicity, we use the notation $P_{ij}^{L}=P_{ij}\otimes I-I\otimes P_{ij}^{t}$ , $P_{ij}=\sigma_{i}\otimes\sigma_{j}$. When $\theta\neq n\pi/2$, this gate is not a Clifford gate. As a starting point, our goal is to create an operation similar to twirling that keeps the functionality of the ideal gate, as in the case of Pauli twirling. If we apply $\mathcal{P}_{xx}=P_{xx}\otimes P_{xx}^{t}$, we get: $\mathcal{P}_{xx}e^{-i\theta\mathcal{H}_{zz}}\mathcal{P}_{xx}=e^{-i\theta\mathcal{H}_{zz}}$ since $\mathcal{P}_{xx}$ and $\mathcal{H}_{zz}$ commute. However, $\mathcal{P}_{xz}$, for example, anti-commutes with $\mathcal{H}_{zz}$, and therefore $\mathcal{P}_{xz}e^{-i\theta\mathcal{H}_{zz}}\mathcal{P}_{xz}=e^{+\imath\theta\mathcal{H}_{zz}}=U_{zz}(-\theta)$. Our PST protocol is based on the fact that if the Pauli operators anti-commute with the Hamiltonian of $U$, we can simply change the sign of the angle (i.e., the sign of the driving fields) to achieve the desired transformation in the end:
$\mathcal{P}_{xz}U_{zz}(-\theta)\mathcal{P}_{xz}=U_{zz}(\theta).$
@@
More generally when the driving Hamiltonian is not a single Pauli we can write $\sum_{\beta}h_{\beta}\mathcal{H}_{P_{\beta}}$ so that $U=e^{-\imath\sum h_{\beta}\mathcal{H}_{P_{\beta}}}$ and then the PST twirling takes the form
$\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}e^{-\imath\mathcal{P}_{\alpha}(\sum_{\beta}h_{\beta}\mathcal{H}_{P_{\beta}})\mathcal{P}_{\alpha}}\mathcal{P}_{\alpha}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}e^{-\imath\sum sgn(\alpha,\beta)h_{\beta}\mathcal{H}_{P_{\beta}}}\mathcal{P}_{\alpha},$ (1)
@@
where $sgn(\alpha,\beta)=tr(P_{\alpha}P_{\beta}P_{\alpha}P_{\beta})/2^{n}$ equals $+1$ if $P_{\alpha}$ and $P_{\beta}$ commute, and $-1$ if they anti-commute. Note that in the definition of $sgn(\alpha,\beta)$ we used Pauli matrices in Hilbert space. Mathematically, expression (1) is straightforward, but it should be considered with execution protocol in mind. The right $\mathcal{P}_{\alpha}$ is executed at the beginning of the gate. Then a modified unitary $U_{\alpha}=e^{-i\sum sgn(\alpha,\beta)h_{\beta}\mathcal{H}_{P_{\beta}}}$ is executed next by changing the sign of some of the control signal. Finally, another $\mathcal{P}_{\alpha}$ is executed at the end of the gate. Importantly, the Hamiltonian of the modified unitary has the same elements as the original Hamiltonian, with only the sign of some terms changing. From an experimental perspective, this implies that implementing PST is no more challenging than implementing the original unitary.
The main operational differences between RC and pseudo twirling are as follows:
@@
  • In RC, the Pauli gates at the beginning and end are different. In pseudo twirling, they are always the same.
@@
  • In RC, the operation between the Pauli gates (the gate itself) is always the same. Conversely, in PST, the operation between the Pauli gates depends on the selected Pauli.
@@
RC and PST are illustrated in Fig. 2a and Fig. 2b respectively. We note that the term ‘pseudo twirling’ reflects not merely an operational difference from standard twirling but also its distinct impact on errors, as discussed in the following two sections. Its effects differs from Pauli twirling in two significant ways: i) it cannot eliminate one type of coherent error, specifically over/under rotations, and ii) it does not transform incoherent errors into a Pauli channel. Nevertheless, we will show and demonstrate that these differences do not diminish the usefulness of pseudo twirling.
I.3 PST Impact on Coherent Errors
While the impact of PST on strong coherent and incoherent errors is should be further studied we do understand its effect on weak error. Through the Magnus expansion [28] and the interaction picture, we find that the PST of a circuit with a coherent error induced by a Hamiltonian term $H_{coh}(t)$ can be expressed as follows:
@@ I.3 PST Impact on Coherent Errors
where $U(t)$ is the ideal evolution operator of the non-twirled gate. $\mathcal{H}_{coh}$ is in Liouville space which is related to the Hilbert space Hamiltonian $H_{coh}$ through the relation
$\mathcal{H}_{coh}=H_{coh}\otimes I-I\otimes H_{coh}^{t}.$ (2)
The first order approximation in $\mathcal{H}_{coh}$ is
@@
$U_{T}+U_{T}\frac{-i}{2^{2n}}\intop U(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}U(t)dt+O(\mathcal{H}_{coh}^{2})$ $=U_{T}+O(\mathcal{H}_{coh}^{2}),$
where the last step follows from the fact that $\frac{1}{2^{n}}\sum_{\alpha=1}^{2^{n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}$ is a Pauli twirling of $\mathcal{H}_{coh}$. As such the twirling eliminates the leading term in the coherent error since $\frac{1}{2^{n}}\sum_{\alpha=1}^{2^{n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}=tr(\mathcal{H}_{coh})=0$. This explains why PST suppresses coherent errors. Higher order terms do not have the structure of (2) and require a more elaborate analysis. Figure 3(a) shows the error map in the Pauli basis of the gate $e^{-i\theta\mathcal{H}_{zz}}$ with an added coherent error. In Figure 3(b) the noise map of the same gate after applying PST is plotted.
Figure 3: Reduction of coherent errors via PST. (a) The noise map of a partial ZZ rotation two-qubit non-Clifford gate $e^{-i\theta\mathcal{H}_{zz}}$ with a coherent error. An ideal gate, devoid of errors, would display as the red pixel in the upper-left corner, while all other elements appear as zeros (white). (b) After applying PST, a significant reduction in noise is observed. However, unlike randomized compiling, the new noise map is not diagonal. Exploring this feature and the PST limitations under stronger coherent errors is part of our research goals.
I.4 PST and Noise
In the context of Pauli twirling, it is evident that the resulting incoherent error exhibits a Pauli error channel structure. However, in the case of PST, this may not hold true, even in the first order. If we revisit the analysis from the previous section but consider a general Lindblad dissipator $\mathcal{L}$ describing a Markovian trace-preserving noise, we obtain:
$U_{T}+U_{T}\frac{-i}{N_{RC}}\intop U(t)^{\dagger}\sum_{i=1}^{2^{2n}}\mathcal{P}_{i}\mathcal{L}\mathcal{P}_{i}U(t)dt+O(\mathcal{L}^{2}),$
@@ I.4 PST and Noise
II PST features - Limitations and Opportunities
II.1 PST and Calibration
@@
There are two types of coherent errors: controllable errors arising from miscalibration of controllable degrees of freedom which we refer to a over/under rotation error, and uncontrollable errors such as crosstalk that do not depend on the control signals (driving fields). While RC makes no distinction and effectively mitigate both types of coherent errors, PST specifically addresses crosstalk and other uncontrollable coherent errors. Crucially over/under rotation errors that are crucial in gate calibration are not affected by PST. This feature opens new possibilities.
@@
Initially, it may seem that this limitation renders PST only a partial solution, as it cannot transform calibration errors into incoherent errors. However this exact limitation can become an advantage in gate calibration. A proper calibration is always a good practice. Even when RC is applicable, miscalibrated gates introduce access noise that makes the QEM task more challenging. PST, on the other hand, does not convert rotation errors into incoherent errors. This characteristic allows for the application of PST during the calibration process, simplifying it by reducing certain coherent errors, such as crosstalk errors, while leaving those associated with the calibration itself untouched. Consequently, integrating PST into the calibration process can facilitate the high-accuracy calibration necessary for circuits with significant depth.
For example consider the case of cross-resonance gate where the leading driving Hamiltonian term is $H_{d}=\alpha\cos\phi\sigma_{z}\otimes\sigma_{x}+\alpha\sin\phi\sigma_{z}\otimes\sigma_{y}$, where $\alpha$ is the amplitude of the drive and $\phi$ is the phase of the drive. After applying PST we have:
$\mathcal{H}_{d}^{PST}$ $=\frac{1}{16}\sum_{\alpha=1}^{16}\frac{tr(P_{\alpha}P_{zx}P_{\alpha}P_{zx})}{4}\mathcal{P}_{\alpha}\mathcal{H}_{d}\mathcal{P}_{\alpha}$
$=\alpha\cos\phi\mathcal{H}_{zx}$
Note that in this case it is not possible to directly control the amplitudes of $\sigma_{z}\otimes\sigma_{x}$ and $\sigma_{z}\otimes\sigma_{y}$ separately. In principle, such control can be achieve by changing the sign of the phase as well. However, by doing the explicit summation find that $\frac{1}{16}\sum_{\alpha=1}^{16}\frac{tr(P_{\alpha}P_{zx}P_{\alpha}P_{zx})}{4}\mathcal{P}_{\alpha}\mathcal{H}_{zy}P_{\alpha}=0$ and therefore the PST summation nulls the $\sigma_{z}\otimes\sigma_{y}$ term even without modifying the sign of the phase. This leads to a simplification in the calibration procedure: rather than fine-tune $\phi$ to zero and then fine tune $\alpha$, it is possible to coarsely set $\phi\sim 0$ and the calibrate $\alpha$ until $\alpha\cos\phi$ achieves the desired value.
II.2 PST application alternatives

### v2

@@ II.2 Introducing the Pseudo Twirling Protocol
To introduce the concept of pseudo twirling we start with the following example. Consider a simple non-Clifford gate, such as the two-qubit gate $\mathcal{U}_{zz}(\theta)=e^{-i\theta\mathcal{H}_{zz}}$, with $\theta\neq k\pi/2$. For simplicity, we use the notation $\mathcal{H}_{zz}=P_{zz}\otimes I-I\otimes P_{zz}^{\intercal}$ , $P_{zz}=\sigma_{z}\otimes\sigma_{z}$. In analogy to Pauli twirling we aim to create an operation that keeps the functionality of the ideal gate. If we apply $\mathcal{P}_{xx}=P_{xx}\otimes P_{xx}^{*}$, we get: $\mathcal{P}_{xx}e^{-i\theta\mathcal{H}_{zz}}\mathcal{P}_{xx}=e^{-i\theta\mathcal{H}_{zz}}$ since $\mathcal{P}_{xx}$ and $\mathcal{H}_{zz}$ commute. However, $\mathcal{P}_{xz}$, for example, anti-commutes with $\mathcal{H}_{zz}$, and therefore $\mathcal{P}_{xz}e^{-i\theta\mathcal{H}_{zz}}\mathcal{P}_{xz}=e^{+\imath\theta\mathcal{H}_{zz}}=\mathcal{U}_{zz}(-\theta)$. Our PST protocol is based on the fact that if the Pauli operators anti-commute with the Hamiltonian of $U$ (or equivalently $\mathcal{U}$ in Liouville space), we can simply change the sign of the angle, i.e., the sign of the driving fields, to achieve the desired transformation in the end:
$\mathcal{P}_{xz}\mathcal{U}_{zz}(-\theta)\mathcal{P}_{xz}=\mathcal{U}_{zz}(\theta).$ (7)
@@
More generally when the driving Hamiltonian is not a single Pauli, we can write $\mathcal{U}=e^{-i\sum h_{\beta}\mathcal{H}_{\beta}}$ and then the PST twirling takes the form
$\mathcal{K}_{PST}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}e^{-i\mathcal{P}_{\alpha}(\sum_{\beta}h_{\beta}\mathcal{H}_{\beta})\mathcal{P}_{\alpha}}\mathcal{P}_{\alpha}=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}e^{-i\sum_{\beta}sgn(\alpha,\beta)h_{\beta}\mathcal{H}_{\beta}}\mathcal{P}_{\alpha},$ (8)
@@
where $sgn(\alpha,\beta)=tr(P_{\alpha}P_{\beta}P_{\alpha}P_{\beta})/2^{n}$ equals $+1$ if $P_{\alpha}$ and $P_{\beta}$ commute, and $-1$ if they anti-commute. Note that in the definition of $sgn(\alpha,\beta)$ we used Pauli matrices in Hilbert space. Mathematically, expression (8) is straightforward, but it should be considered with execution protocol in mind. The right $\mathcal{P}_{\alpha}$ is executed at the beginning of the gate. Then a modified unitary $\mathcal{U}_{\alpha}=e^{-i\sum_{\beta}sgn(\alpha,\beta)h_{\beta}\mathcal{H}_{\beta}}$ is executed next by changing the sign of some of the control signals. Finally, another $\mathcal{P}_{\alpha}$ is executed at the end of the gate. Importantly, the Hamiltonian of the modified unitary has the same elements as the original Hamiltonian, with only the sign of some terms changing. Thus, under the assumption that the signs of the driving Hamiltonian terms are controllable, the PST implementation is no more challenging than implementing the original unitary.
The main operational differences between RC and pseudo twirling are:
@@
  • In RC, the Pauli gates at the beginning and at the end are different. In pseudo twirling, they are always the same.
@@
  • In RC, the operation between the Pauli gates (the original gate) is always the same. Conversely, in PST, the operation between the Pauli gates depends on the selected Pauli.
@@
RC and PST are illustrated in Fig. 2a and Fig. 2b respectively. We note that the term ‘pseudo twirling’ reflects not merely an operational difference from standard twirling but also its distinct impact on errors, as discussed in the following two sections. Its effects differs from Pauli twirling in two significant ways: i) it cannot eliminate over/under rotations errors, and ii) in general, it does not transform incoherent errors into a Pauli channel. Nevertheless, we show and demonstrate that these differences do not diminish the usefulness of pseudo twirling.
II.3 PST Impact on Coherent Errors
In this section we study the impact of PST on coherent errors. Through the Magnus expansion [34] in the interaction picture, we find that the application of PST in a circuit with a coherent error induced by a Hamiltonian term $\mathcal{H}_{coh}(t)$ leads to the following evolution operator
@@ II.3 PST Impact on Coherent Errors
where $\mathcal{U}(t)$ is the ideal evolution operator of the non-twirled gate and $\mathcal{U}_{T}=\mathcal{U}(T)$. The second and higher order terms in the Magnus expansion have been neglected. $\mathcal{H}_{coh}$ is in Liouville space which is related to the Hilbert space Hamiltonian $H_{coh}$ through Eq. (2). The first order Taylor expansion in $\mathcal{H}_{coh}$ is
@@
$\mathcal{U}_{T}+\mathcal{U}_{T}\frac{-i}{2^{2n}}\intop_{0}^{T}\mathcal{U}(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}\mathcal{U}(t)dt+O(\mathcal{H}_{coh}^{2})$ $=\mathcal{U}_{T}+O(\mathcal{H}_{coh}^{2}),$ (10)
where the last step follows from the fact that $\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}$ is a Pauli twirling of $\mathcal{H}_{coh}$. As such, the twirling eliminates the leading term in the coherent error since $\sum_{\alpha=1}^{2^{2n}}P_{\alpha}H_{coh}P_{\alpha}=tr(H_{coh})I$ and $\sum_{\alpha=1}^{2^{2n}}P_{\alpha}^{*}H_{coh}^{\intercal}P_{\alpha}^{*}=tr(H_{coh})^{*}I=tr(H_{coh})I$ so that the two terms in $\mathcal{H}_{coh}$ cancel each other. This explains why PST suppresses coherent errors.
Next, we study the effect of PST on the second order Taylor approximation in $\mathcal{H}_{coh}$. This term is important in light of the fact the first order is eliminated by PST. Furthermore, this analysis will unravel the structure of the resulting noise Channel. The second order Taylor expansion of (9) leads to following 2nd order term in $\mathcal{K}_{pst}$
[…]
$-(H_{eff,\alpha}\otimes I-I\otimes H_{eff,\alpha}^{\intercal})^{2}$ $=2H_{eff,\alpha}\otimes H_{eff,\alpha}^{\intercal}-H_{eff,\alpha}^{2}\otimes I-I\otimes(H_{eff,\alpha}^{2})^{\intercal}.$ (12)
Since $H_{eff,\alpha}$ is Hermitian this expression has a Lindblad form $\mathcal{L}(A)=A\otimes A^{*}-\frac{1}{2}(A^{\dagger}A)\otimes I-\frac{1}{2}I\otimes(A^{\dagger}A)^{\intercal}$ with a Hermitian dissipator $A=H_{eff,\alpha}$. We conclude that PST nulls the first order of the coherent error and the second order produces a Hermitian Lindbladian. That is
$\mathcal{K}_{PST}$ $\simeq\mathcal{U}_{T}e^{\mathcal{L}_{eff}},$ (13)
$\mathcal{L}_{eff}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{L}(H_{eff,\alpha}).$ (14)
Since $\mathcal{L}_{eff}=\mathcal{L}_{eff}^{\dagger}$ the resulting noise channel $\mathcal{N}_{PST}=\mathcal{U}_{T}^{\dagger}\mathcal{K}_{PST}\simeq e^{\mathcal{L}_{eff}}$ is also Hermitian up to higher order corrections. We emphasize that unlike RC, $\mathcal{N}_{PST}$ is not a Pauli channel although it is approximately Hermitian. This is easily observed when plotting $\mathcal{N}_{PST}$ in the Pauli basis. The noise map is not diagonal as in RC. This fundamental difference highlights that PST is not a variant of RC or a specific way of implementing RC. It is a different protocol with different analytical features and a different operating regime.
If $H_{coh}$ is a Pauli then $\mathcal{P}_{\alpha}\mathcal{H}_{coh}\mathcal{P}_{\alpha}=\pm\mathcal{H}_{coh}$. However since $\mathcal{L}(A)=\mathcal{L}(-A)$ we get that $\mathcal{L}_{eff}$ simplifies to
[…]
We leave the study of higher order terms to future research.
II.4 PST and Noise
In the context of Pauli twirling, it is evident that the resulting incoherent error exhibits a Pauli error channel structure. However, in the case of PST, this may not hold true, even in the first order. If we repeat the analysis from the previous section but consider a general Lindblad dissipator $\mathcal{L}$ describing a Markovian trace-preserving noise, we obtain
$\mathcal{U}_{T}+\mathcal{U}_{T}\frac{-i}{2^{2n}}\intop\mathcal{U}(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{L}\mathcal{P}_{\alpha}\mathcal{U}(t)dt+O(\mathcal{L}^{2}),$ (16)
@@ II.4 PST and Noise
III PST features - limitations and opportunities
III.1 PST, Additional Types of Coherent Errors, and Calibration
@@
Coherent errors comprise of either controllable errors arising from miscalibration of controllable degrees of freedom, or uncontrollable errors such as crosstalk that do not depend on the control signals (driving fields). While RC makes no distinction and effectively mitigate both types of coherent errors, PST addresses all coherent error excluding controllable rotation errors (over or under rotation) which play a key role in gate calibration. This feature opens new possibilities.
@@
At first glance, it may seem that this limitation renders PST only a partial solution, as it cannot transform controllable rotation errors into incoherent errors. However, this exact limitation can become an advantage in gate calibration. A proper calibration is always a good practice. When RC is applicable, it will convert the miscalibration error into excessive noise that makes the QEM task more challenging. PST, on the other hand, does not convert controllable rotation errors into incoherent errors. This characteristic allows for the application of PST during the calibration process. By reducing all other coherent errors, and leaving only those associated with the calibration itself, it simplifies the calibration process and improves its accuracy. The high accuracy calibration that PST facilitates, is imperative for circuits with significant depth.
As a concrete scenario we consider a simplified cross-resonance Hamiltonian $H_{d}=a\cos\phi\,\sigma_{z}\otimes\sigma_{x}+a\sin\phi\,\sigma_{z}\otimes\sigma_{y}+\zeta\sigma_{z}\otimes\sigma_{z}$, where $a$ is the amplitude of the drive, $\phi$ is the phase of the drive, and $\zeta$ is the crosstalk coupling between the qubits. See [29] for additional terms. Since $\sigma_{z}\otimes\sigma_{x}$ is the target Hamiltonian, a deviation of the coefficient $a\cos\phi$ from the desired value is considered as a rotation error. Since it is possible to change the sign of this coefficient we consider it as a controllable rotation error. The term $a\sin\phi\,\sigma_{z}\otimes\sigma_{y}$ is a controllable non-rotational error and the last crosstalk term is an uncontrollable non-rotational error. Such scenarios motivate a more formal and accurate classification of coherent errors. Let us assume we have a Hamiltonian of the form
$\mathcal{H}=f(a)\mathcal{H}_{\beta}+\sum_{\gamma\neq\beta}g_{\gamma}(a)\mathcal{H}_{\gamma},$ (17)
where $\mathcal{H}_{\beta}$ is the target Pauli Hamiltonian (multi-qubit Pauli), $\mathcal{H}_{\gamma}$ represents Pauli coherent error terms, and $a$ is the parameter being controlled. Without loss of generality we assume $\mathcal{H}_{\gamma}\neq\mathcal{H}_{\beta}$. If they are equal it is possible to absorb $g(a)$ in the definition of $f(a)$. For brevity, in what follows we consider a single $\mathcal{H}_{\gamma}$ term. Next, we decompose $f$ and $g$ into even and odd contributions, i.e. $f(a)=f_{e}(a)+f_{o}(a)$ where $f_{e}(a)=\frac{1}{2}[f(a)+f(-a)]$ and $f_{o}(a)=\frac{1}{2}[f(a)-f(-a)]$. We refer to the odd terms as controllable since changing the sign of $a$ changes the sign of the terms in the Hamiltonian as needed for PST. Similarly, the even functions represent uncontrollable degrees of freedom from the point of view of PST. Miscalibration of $f(a)$, regardless if the odd or even term is the problem, leads to rotation errors. $g(a)$ stands for non-rotation errors that stems, for example, from crosstalk or parasitic terms in the control signals that generate the target term $\mathcal{H}_{\beta}$. In what follows, it is shown that PST mitigates the terms proportional to $f_{e}(a),g_{o}(a),$ and $g_{e}(a)$, while $f_{o}(a)$ remains unaffected.
While $f_{e}(a),g_{o}(a),$ and $g_{e}(a)$ are small since they represent coherent errors, $f_{o}(a)$ may be potentially large. This is addressed by moving to the interaction picture of $f_{o}(a)\mathcal{H}_{\beta}$. After applying PST using $\mathcal{P_{\alpha}}$ and $a\to sgn(\alpha,\beta)a$, the first order expansion [as in (10)], yields
$\mathcal{K}_{PST}$ $=\mathcal{U}_{T}^{o}+\mathcal{U}_{T}^{o}\frac{-i}{2^{2n}}\intop_{0}^{T}\mathcal{U}^{o}(t)^{\dagger}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\{f_{e}(a)\mathcal{H}_{\beta}+[g_{e}(a)+sgn(\alpha,\beta)g_{o}(a)]\mathcal{H}_{\gamma}\}\mathcal{P}_{\alpha}\mathcal{U}^{o}(t)dt$
$+O(f_{e}^{2},g_{e}^{2},g_{o}^{2},g_{e}g_{o},f_{e}g_{o},f_{e}g_{e}),$
where we have used the parities of $f_{e}$, $g_{e}$ and $g_{o}$ to simplify the expression. $\mathcal{U}^{o}(t)$ is the evolution operator generated by $f_{o}(a)\mathcal{H}_{\beta}$ ($\mathcal{U}_{T}^{o}=\mathcal{U}^{o}(T))$. $\mathcal{U}_{T}^{o}$ contains all potential controlled rotation errors. We can now focus on the effective Hamiltonian:
$\mathcal{H}_{eff}$ $=\frac{1}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\{f_{e}(a)\mathcal{H}_{\beta}+[g_{e}(a)+sgn(\alpha,\beta)g_{o}(a)]\mathcal{H}_{\gamma}\}\mathcal{P}_{\alpha}$ (18)
$=\frac{g_{o}(a)}{2^{2n}}\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}_{\alpha},$ (19)
where we have used the fact that $f_{e}(a)$ and $g_{e}(a)$ do not depend on $\alpha$ and therefore the regular twirling property $\sum_{\alpha=1}^{2^{2n}}\mathcal{P}_{\alpha}\mathcal{H}_{\beta(\gamma)}\mathcal{P}_{\alpha}=0$ nulls these terms. In Appendix II, we show that when $\mathcal{H}_{\gamma}\neq\mathcal{H}_{\beta}$ it holds that $\sum_{\alpha=1}^{2^{2n}}sgn(\alpha,\beta)\mathcal{P}_{\alpha}\mathcal{H}_{\gamma}\mathcal{P}_{\alpha}=0$ and we get $\mathcal{H}_{eff}=0$ so that
$\mathcal{K}_{PST}=\mathcal{U}_{T}^{o}+O(f_{e}^{2},g_{e}^{2},g_{o}^{2},g_{e}g_{o},f_{e}g_{o},f_{e}g_{e}).$ (20)
In the cross-resonance case discussed above $f_{o,zx}(a)=a\cos\phi$, $g_{o,zx}(a)=a\sin\phi$, $f_{e,zx}=0,g_{e,zx}=0$,$g_{o,zz}=0$ and $g_{e,zz}=\zeta$. After PST the effective driving Hamiltonian is
$\mathcal{H}_{d}^{PST}$ $=a\cos\phi\mathcal{H}_{zx}.$ (21)
This leads to a simplification in the calibration procedure: rather than fine-tune $\phi$ to zero and then fine-tune $a$ (e.g. to $\pi/4$), it is possible to coarsely set $\phi\sim 0$ and then calibrate $a$ until $a\cos\phi$ achieves the desired value. Furthermore, PST also suppresses crosstalk to neighboring qubits which improves the accuracy of calibration.
III.2 Alternative Ways for Applying PST in a Circuit